OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line. Express your answer in the form y=mx+b (4,9); 7x+y=4

OpenStudy (dumbcow):

y = 1/7x +b 9 = 1/7(4) +b b = 59/7 y = 1/7x +59/7

OpenStudy (anonymous):

the slope of the line 7x+y=4 is -7. so the slope of the perpendicular line will be 1/7.(as -7*(1/7)=-1). so the line be y=x/7+c line passes through (4,9) 9=4/7+c=>c=59/7 so the line is y=x/7+59/7